OpenStudy (anonymous):

Solve the equation by completing the square. 3x2 + 18x = -15

OpenStudy (anonymous):

3x^2+18x+15=0 (3x+3)(x+5)=0 3x+3=0 3x=-3 x=-1 or x+5=0 x=-5

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (primeralph):

@nique7996 Your answer and method are both wrong.

OpenStudy (primeralph):

@marsss That is not completing the square method.

OpenStudy (primeralph):